OpenStudy (anonymous):

If the vertices of a triangle are A(5, -1), B(-3, 1), and C(1, 5), what are the lengths of its medians? (Hint: Review the definition of a median of a triangle and the distance formula.) Median AB = Median BC = ? sqrt ? Median CA = sqrt ?

OpenStudy (anonymous):

you need to get all the midpoints of the three sides and then calculate the distance between each two midpoints and by that u'll have the length of all the medians

OpenStudy (anonymous):

to find the coordinates of a midpoint here's the rule : x=x1+x2/2... y =y1+y2/2 to find the distance between any two points here's the rule : \[\sqrt{(x2-x1)^2 +(y2-y1)^2}\]